Average Rate of Return, Cash Payback Period, Net Present Value Method for a Service Company

Spanish Peaks Railroad Inc. is considering acquiring equipment at a cost of $344,000. The equipment has an estimated life of 10 years and no residual value. It is expected to provide yearly net cash flows of $43,000. The company's minimum desired rate of return for net present value analysis is 15%.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

Compute the following:

a. The average rate of return, giving effect to straight-line depreciation on the investment. If required, round your answer to one decimal place.
%

b. The cash payback period.
8 years

c. The net present value. Use the above table of the present value of an annuity of $1. Round to the nearest dollar. If required, use a minus sign to indicate negative net present value for current grading purpose.

Present value of annual net cash flows $
Amount to be invested $
Net present value $
Answer #1

a. Annual depreciation = (344000-0)/10 = 34400

Net income = 43000+34400 = $77400

Average rate of return = Net income/Investment

= 77400/344000

= 22.5%

b. Cash payback period = 344000/43000 = 8 years

c.

Present value of annual net cash flows (43000*5.019) $215817
Amount to be invested $344000
Net present value $-128183